A ball is shot from the ground into the air. After 20s it’s velocity is v=(9.6i+6.5j) m/s, with i horizontal and j upwards. To what maximum height does the ball rise? What are the magnitude of the ball’s velocity just before it hits the ground?


1
Expert's answer
2021-03-23T11:10:55-0400
"v_{y0}=6.5+9.8(20)=202.5\frac{m}{s}\\v^2=2gh\\202.5^2+9.6^2=2(9.8)h\\h=2100\ m"